OpenStudy (anonymous):

how to solve this? 5x²-15x=2(3-x)

OpenStudy (anonymous):

Factorise the left side first by taking out the greatest common factor, what do you get?

OpenStudy (harsimran_hs4):

just remember everytime you cut out the common factors from both the sides, put those expressions as zero also otherwise you may loose out some of the factors

OpenStudy (anonymous):

5x^2-15x=6-2x add all factors to one side 5x^2-13x -6=0 (5x^2-15x)+(2x-6) factor 5x(x-3)+2(x-3) (5x+2)+(x-3) then set them equal to zero (5x+2)=-2/5 (x-3)=3

OpenStudy (harsimran_hs4):

yes ! they are factors so no + sign in 2nd last step